OpenStudy (anonymous):

find the values of x where the extreme values of the function y=x^3-7x-5x occur.

OpenStudy (anonymous):

(a) x=-0.33, x=5 (b) x=0.33, x=-5 (c) x=0.33, x=5 (d) x=2, x=5 (e) x=0.65 ,x=7.65